If the probability of picking a red marble out of a bag is 0.4 and the probability of choosing a green marble out of the same bag is 0.25 what is the probability of randomly choosing a red or green marble out of the bag with one pick
So...
p of red=0.4
p of green=0.25
You are looking for the p of red or green
when we have an "or" we add probability so....
.4+.25=0.65